This problem states, log base 9 (x-7) + log base 9 (x-7) is equal to 1.
I have done this over and over and can not seem to come up with the right answer. :)
log( 9, (x-7) + log( 9, (x-7) = 1

Answer by nerdybill(7384)

Applying "log rules":
log9(x-7) + log9(x-7) = 1
log9((x-7)(x-7)) = 1
(x-7)(x-7) = 9^1
(x-7)(x-7) = 9
x^2-7x-7x+49 = 9
x^2-14x+49 = 9
x^2-14x+40 = 0
(x-10)(x-4) = 0
.
x = {10,4}